Mr. A and Mr. B have the following respective demands


Pa= 25- 5X

Pb= 11- X

a. Graph them on separate graphs

b. Mr A has no X's while Mr. B has 10 X's. What are the marginal and total values at this quantity? Using either method discussed in the chapter, determine how much each will have after trade? What will be their marginal values at this quantity?


1
Expert's answer
2022-04-19T09:59:07-0400

"P_a=25-5X,"

"P_b=11-X,"

"P_a=P_b,"

"25-5X=11-X,"

"14=4X,"

"X=3.5,"

"P_a=P_b=7.5."